a dam is 12.5m deep if it hold a liquid of density 1250kgm find the pressure at the base.[g=10m/s]
Answer by fractalier(6550)

If I am not mistaken, hydrostatic pressure of a vertical liquid is given by
P = rho*g*h = (1250 kg/m^3)(10 m/s^2)(12.5 m) = 156.25 kPa